OpenStudy (anonymous):

Using a trig identity, write (equation in question) using only one cosine function

OpenStudy (anonymous):

x(t)=5cos(4t)-2sin(4t)

OpenStudy (anonymous):

Use the identity \[\cos(\alpha-\beta)=\cos\alpha\cos\beta+\sin\alpha\sin\beta\]

OpenStudy (anonymous):

so, the alpha and betas are both 4t right? What do I do with the integers before the trig functions?

OpenStudy (anonymous):

If both alpha and beta were 4t, you'd have \[\cos0=\cos^24t+\sin^24t,\] which, although true, doesn't help you. I'll try to work it out by hand for myself; I haven't done this particular problem in a while.

OpenStudy (anonymous):

oh wow right, so I must have to split up the (4t) into 2 components somehow, to get and alpha and beta, and then find the difference between them.....

OpenStudy (anonymous):

i found a trig identity that should help, but i don't know how to get it into one cosine function. \[Acos(wt)+Bsin(wt)=Ccos(wt-\gamma)\]

OpenStudy (anonymous):

where \[C=\sqrt(A^2+B^2)\] sin(gamma)=B/C cos(gamma)=A/C

OpenStudy (anonymous):

Starting with the identity, and letting k be some constant: \[k\cos(\alpha-\beta)=k\cos\alpha\cos\beta+k\sin\alpha\sin\beta\] You want the right side to look like the given expression for x(t), so we'll equate the right side to it: \[k\cos\alpha\cos\beta+k\sin\alpha\sin\beta=5\cos(4t)-2\sin(4t)\] Letting alpha = 4t, \[k\cos\beta\cos(4t)+k\sin\beta\sin(4t)=5\cos(4t)-2\sin(4t)\] Matching up the coefficients, you have \[\begin{cases}k\cos\beta=5\\ k\sin\beta=-2\end{cases},\text{ or equivalently, }\begin{cases}\cos\beta=\frac{5}{k}\\ \sin\beta=\frac{-2}{k}\end{cases}\] Draw up a reference triangle that satisfies the system:  Solving for k yields \[k^2=5^2+(-2)^2\Rightarrow k=\sqrt{29}\] So, you're left with \[\begin{align*}k\cos(\alpha-\beta)&=5\cos(4t)-2\sin(4t)\\ k\cos\beta\cos(4t)+k\sin\beta\sin(4t)&=\\ \sqrt{29}\cos(4t-\beta)&=\end{align*}\] From the triangle you know that \[\tan\beta=\frac{-2}{5},\text{ so }\beta=\tan^{-1}\frac{-2}{5}\] So, you get \[x(t)=\sqrt{29}\cos\left( 4t-\tan^{-1}\left(-\frac{2}{5}\right)\right)\]

OpenStudy (anonymous):

i got down to \[\sqrt(29)*\cos(4t-\cos^{-1} (-2/\sqrt29)\] but the question wants the answer in terms of only one cosine function

OpenStudy (anonymous):

and unfortunately the same holds true for your answer :(

OpenStudy (anonymous):

It's still technically in terms of cosine. The inverse tangent term in the result is just a number.

OpenStudy (anonymous):

woah ok so i put in your answer any way and it worked....thanks you so much

OpenStudy (anonymous):

tan(some number) would also be some number. tan(t) is a function of t.

OpenStudy (anonymous):

ok I get it! thank you.....how do you give medals cuz you deserve 2 haha